# Efficient Camera Exposure Control for Visual Odometry via Deep Reinforcement Learning

### v1

@@ Abstract
The stability of visual odometry (VO) systems is undermined by degraded image quality, especially in environments with significant illumination changes. This study employs a deep reinforcement learning (DRL) framework to train agents for exposure control, aiming to enhance imaging performance in challenging conditions. A lightweight image simulator is developed to facilitate the training process, enabling the diversification of image exposure and sequence trajectory. This setup enables completely offline training, eliminating the need for direct interaction with camera hardware and the real environments. Different levels of reward functions are crafted to enhance the VO systems, equipping the DRL agents with varying intelligence. Extensive experiments have shown that our exposure control agents achieve superior efficiency—with an average inference duration of 1.58 ms per frame on a CPU—and respond more quickly than traditional feedback control schemes. By choosing an appropriate reward function, agents acquire an intelligent understanding of motion trends and anticipate future illumination changes. This predictive capability allows VO systems to deliver more stable and precise odometry results. The codes and datasets are available at https://github.com/ShuyangUni/drl_exposure_ctrl.
@@ I Introduction
Current camera exposure control methods for VO tasks are principally categorized into gradient-based and function-fitting approaches. Gradient-based methods employ a differentiable image metric to guide exposure optimization and use a feedback control mechanism to iteratively refine exposure settings, ensuring stability and cost-effectiveness. However, these methods are limited to the requirement of differentiable metrics and typically require multiple frames to achieve optimal exposure due to the inherent lag of the feedback control system. In contrast, function-fitting methods establish a relationship between the image metric and exposure, typically offering a one-shot solution that responds instantly and can accommodate any metric design, albeit at the cost of significant computational resources.
@@
Existing exposure control methods often struggle under challenging conditions. Transitions through environments with drastically differing illumination, such as tunnels [1], can destabilize the frontend of VO systems: Gradient-based methods produce over-saturated images due to delays in parameter updates due to their feedback mechanisms. Function-fitting methods, despite their rapid response, result in significant disparities between adjacent frames. Both methods may ultimately lead to the failure of interframe tracking. Developing a robust camera exposure control method that can intelligently manage such extreme conditions will markedly improve the performance of VO systems.
In this study, we leverage a deep reinforcement learning (DRL) framework to enhance the system’s intelligence and robustness in extreme lighting conditions. DRL integrates high-level information directly into its reward design and efficient real-time control outputs. As a data-driven approach, DRL assimilates assimilates specific task and scenario requirements from data, enabling intelligent and adaptive responses. Several existing methods [1, 2] have pursued similar avenues. However, these methods necessitate online interactions during the training phase, relying on specially designed devices to interact with real-world environments, which introduces challenges in implementation and sample efficiency—a common limitation in DRL applications. The contributions of our work include:
@@ II-A Optimization-based Exposure Control
Gradient-based methods [3, 4, 5] compute the gradient of the image metric relative to exposure to determine the direction for parameter updates. These methods employ iterative optimization to refine exposure settings progressively. Shim et al. [3] applied $\gamma$-correction to simulate exposure variations, assessing image gradient magnitudes, The target exposure is relatedly determined by the gamma value derived by a fifth-order polynomial. Zhang et al. [4] enhanced Shim’s metric and rigorously validated their derivatives. Han et al. [5] combined gradient-based iterative search with image simulations, integrating optical flow to evaluate camera self-motion, thus enabling adjustments to gain and exposure time to mitigate motion blur. These methods typically exhibit reactive behavior, adjusting only in response to changes in scene conditions, which can lead to over-saturation in dynamically lit environments.
Function-fitting methods [6, 7] utilize synthetic imaging techniques to generate images at varied exposure levels, targeting global optimality directly. Kim et al. [6] developed a mixed image quality metric and produced synthetic images from a fixed exposure seed image, with Bayesian optimization to simultaneously refine exposure time and gain settings. Zhang et al. [7] integrated image bracketing patterns with synthetic techniques to periodically survey a high dynamic range (HDR) spectrum of the scene, thereby improving the capture of scene irradiance. Function-fitting approaches provide rapid responses, potentially offering optimal settings from a single image capture; however, they require significant computational resources due to the exhaustive global optimization process.
@@ II-B Reinforcement Learning-based Exposure Control
Recent advancements in deep reinforcement learning has markedly broadened their applications in real-world robotic tasks, including locomotion control [8, 9], perception [10, 11], and decision making [12, 13].
Existing DRL-based exposure control methods [1, 2] suffer the online training scheme, which is inefficient and in the risk of non-convergence. Tomasi et al. [1] utilized a deep convolutional neural network (CNN) to dynamically adjust camera gain and exposure time settings. Faced with the challenge of directly labeling optimal exposure parameters, they devised a self-supervised labeling technique using dual cameras and a specialized capture pattern. However, their approach of online data collection combined with offline agent training presents complexities and inefficiencies in implementation. Additionally, the infrequent policy updates within their system may hinder the generalization capabilities of the agent. Lee et al. [2] proposed a device-specific method to facilitate online training. They vectorized input images to eliminate CNN modules, thereby reducing model complexity. They introduced a progressive lighting curriculum that incrementally exposed the agent to increasingly complex training scenarios, which streamlined the training process. Their lightweight agent exhibited improved performance across various visual tasks. Nonetheless, the dependence on a specialized stable device limits their method’s generalizability and fails to address camera motion. Moreover, their approach to image vectorization sacrifices essential high-level visual information, such as feature detection and tracking.
@@ III-B Our System Design
In this work, we employ DRL to train agents for camera exposure control. Existing research [1, 2] directly outputs the dual control parameters from the agent network, requiring a fully online interactive scheme, which is both time-consuming and inefficient. We split the exposure control problem into two distinct modules. The first module, referred as $\mathbf{ModuleExpo}$, determines the optimal exposure $e$ for the coming image, formulated as
@@ IV-A Photometric Synthesis Process
Further details can be found in our prior work [7].
Inspired from Gamache et al. [17], we leverage bracketing image techniques to encompass a wide range of irradiances present within a scene. Specifically, we employ a bracketing capture consisting of $5$ images (Fig. 3(a)) with exposure times of $50$ us, $200$ us, $1$ ms, $5$ ms, and $20$ ms. Our previous research [7] indicates that the synthesis process encounters constraints when transitioning from long to short exposures, as data in over-exposed regions becomes irretrievable. Consequently, when an agent requires an image with a specific target exposure, we select the seed image whose exposure time is just below the target, and then serves it as the basis for synthesizing, thus enabling effective interaction.
@@ IV-B Motion Augmentation
We realize the necessity of incorporating a broad spectrum of motion dynamics within our training datasets. However, once the training trajectories are collected, they remain invariant. Additionally, employing a bracketing technique to widen the range of irradiance inherently diminishes the simulator’s frame rate output. Specifically, using a bracket of $5$ reduces the frame rate to one-fifth of the original sequence rate. This reduction in frame rates indirectly affects the perceived speed of camera movement, thus introducing discrepancies between training and inference conditions.
@@ V-C3 Pose Error Reward
Sequence Built-in Shim Zhang DRL-stat DRL-feat DRL-pose Dura. (s) Vel. (m / s)
Courtyard $\times$1 1.42 / 0.29 2.02 / 0.31 2.55 / 0.31 2.17 / 0.31 1.77 / 0.33 2.04 / 0.35 128.80 1.40
$\times$2 - (1) 7.98 / 0.88 1.64 / 0.49 1.50 / 0.49 1.75 / 0.52 2.38 / 0.58 64.40 2.80
Corridor $\times$1 1.98 / 0.22 2.24 / 0.24 2.19 / 0.24 1.96 / 0.23 0.82 / 0.21 3.00 / 0.32 173.50 0.93
$\times$2 4.28 / 0.39 (1) 5.26 / 0.44 (1) 4.88 / 0.45 (1) 4.13 / 0.43 (1) 1.29 / 0.32 7.48 / 0.64 (2) 86.75 1.85
Parking $\times$1 2.40 / 0.43 3.14 / 0.43 1.44 / 0.42 2.34 / 0.42 1.98 / 0.41 2.69 / 0.51 128.00 1.71
$\times$2 - (1) - (1) - (2) - (3) 1.84 / 0.62 - (1) 64.00 3.41
Switch $\times$1 2.52 / 0.34 (1) 3.38 / 0.35 (1) 1.81 / 0.27 (1) 2.30 / 0.32 (1) 2.67 / 0.34 (1) 1.65 / 0.23 40.79 1.11
@@ VI-B Experiment: Visual Odometry
Several methods are implemented as the baseline comparison. Built-in is a feedback control method from FLIR cameras, aimed at optimizing average image brightness to a moderate value. Shim and Zhang are two methods target on optimizing image gradient magnitude, but with feedback control and function fitting schemes, respectively. Since we designed the three reward functions, we trained DRL-stat, DRL-feat, and DRL-pose for comparison.
(a) Countyard Case.
(b) Parking Case.
(c) Switch Case.
Figure 5: Illustrations of different scenarios in our evaluation datasets.
Quantitative results are shown in Table. II. In the Courtyard sequence, all methods perform well expect for Built-in at double speed, which failed to adjust quickly enough to rapid motion (Fig. 5(a)), leading to over-exposure and tracking failures. In the sequences of Corridor and Parking with drastic lighting variations, DRL-feat shows the best performance, maintaining interframe tracking even under double speed (Fig. 5(b)). In the Switch sequence, characterized by unpredictable lighting changes of $4$ times, all methods were interrupted by the most dramatic transition from dark to light, expect DRL-pose, shown in Fig.5(c). This is because DRL-pose prefers a lower exposure: DRL-pose aims to directly minimize inter-frame pose errors. This enhances the agent’s preference for stable structural features over unstable textural features, which are plentiful but less reliable. Low exposure settings help obscure textural features and prevent blooming effects around structural features, which are typically delineated by contrasts between light and dark areas.
Additionally, we conducted a static experiment in the Switch scene to compare the response speeds of different methods to abrupt unpredictable lighting changes. The response performance is illustrated in Fig. 6. The function fitting method from Zhang exhibited the quickest response. Our DRL-based methods, equivalent to the mechanism of feedback control, all respond faster than the traditional feedback control methods.
Figure 6: The experiment of the reaction speed with a switch case of light off then on. The method of Zhang performs the fastest because of its principle of one step control. All our DRL methods (DRL-stat, DRL-feat, and DRL-pose) respond faster than feedback control methods (Built-in and Shim).
@@ VI-D Experiment: Time Consumption
Our training process does not necessitate any interaction with camera hardware or the real environment, rendering it significantly faster than previous approaches [1, 2]. The training durations for our three agents—DRL-stat, DRL-feat, and DRL-pose—were $3$ hours $13$ minutes, $9$ hours $35$ minutes, and $10$ hours $33$ minutes, respectively, all trained for $10000$ episodes. The average processing time per frame was recorded at $2.3$ ms, $6.9$ ms, and $7.6$ ms, respectively.

### v2

@@ Abstract
The stability of visual odometry (VO) systems is undermined by degraded image quality, especially in environments with significant illumination changes. This study employs a deep reinforcement learning (DRL) framework to train agents for exposure control, aiming to enhance imaging performance in challenging conditions. A lightweight image simulator is developed to facilitate the training process, enabling the diversification of image exposure and sequence trajectory. This setup enables completely offline training, eliminating the need for direct interaction with camera hardware and the real environments. Different levels of reward functions are crafted to enhance the VO systems, equipping the DRL agents with varying intelligence. Extensive experiments have shown that our exposure control agents achieve superior efficiency—with an average inference duration of 1.58 ms per frame on a CPU—and respond more quickly than traditional feedback control schemes. By choosing an appropriate reward function, agents acquire an intelligent understanding of motion trends and can anticipate future changes in illumination. This predictive capability allows VO systems to deliver more stable and precise odometry results. The code and dataset are open source on https://github.com/ShuyangUni/drl_exposure_ctrl.
@@ I Introduction
Traditional exposure control methods are divided into gradient-based [1, 2, 3] and function-fitting approches [4, 5], each can attain satisfactory outcomes under standard conditions. Nevertheless, these methods manifest limitations in environments with drastical illumination changes, such as swift transitions through tunnels [6]. The gradient-based methods suffer from restricted feedback control speed, resulting in the inability to update exposure parameters with rapid scene alterations, consequently leading to overexposure or underexposure. The function-fitting approaches, formulating a correlation between image metrics and exposure, provide an instantaneous, one-shot solution adaptable to various metric designs. However, the substantial computational demand leads to infrequent parameter updates. While this frequency limitation may prevent underexposure, abrupt parameter adjustments can introduce significant variability in image content, thereby undermining feature tracking efficacy.
@@
Several researchers [6, 7] have incorporated deep reinforcement learning (DRL) into exposure control tasks, yielding numerous benefits. DRL-based systems employ an end-to-end framework from image inputs to control output, obviating the need for complex system design. the reward-based nature of reinforcement learning facilitates the integration of direct optimization targets, such as the number of features. These targets, which are non-differentiable and labor-intensive to compute using traditional techniques, are readily accommodated within the DRL framework. However, current DRL-based methods that rely on online interaction face multiple challenges. The design of the requisite custom hardware is complex and is difficult for the comunity to reimplement. Additionally, the mode of online interaction leads to inefficient data sampling, where each data sequence is employed only once and becomes obsolete after a policy update. This inefficiency in data utilization hampers the system’s ability to adequately update the policy within the constraints of limited training periods. Therefore, to enhance training stability, it becomes necessary to integrate specific mechanisms [6] or employ curriculum learning [7].
In this study, we also propose a DRL-based exposure control method, focusing on the intelligence and robustness of the VO system under complex lighting conditions. We decouple data collection from model training by adopting an offline training approach to improve sample efficiency. Compared to prior methods, our framework demonstrates enhanced convergence efficiency and simplicity. The contributions of our work include:
@@ II-A Optimization-based Exposure Control
Gradient-based methods [1, 2, 3] compute the gradient of the image metric relative to exposure to determine the direction for parameter updates. These methods employ iterative optimization to refine exposure settings progressively. Shim et al. [1] applied $\gamma$-correction to simulate exposure variations, assessing image gradient magnitudes. Zhang et al. [2] enhanced Shim’s metric and rigorously validated their derivatives. Han et al. [3] integrated optical flow to evaluate camera self-motion, thus enabling adjustments to gain and exposure time to mitigate motion blur. These methods typically exhibit reactive behavior, adjusting only in response to changes in scene conditions, which can lead to over-saturation in dynamically lit environments.
Function-fitting methods [4, 5] utilize synthetic imaging techniques to generate images at varied exposure levels, targeting global optimality directly. Kim et al. [4] developed a mixed image quality metric and produced synthetic images from a fixed exposure seed image, with Bayesian optimization to simultaneously refine exposure time and gain settings. Zhang et al. [5] integrated image bracketing patterns with synthetic techniques to periodically survey a high dynamic range (HDR) spectrum of the scene. Function-fitting approaches provide rapid responses, potentially offering optimal settings from a single image capture; however, they require significant computational resources due to the exhaustive global optimization process.
@@ II-B Reinforcement Learning-based Exposure Control
Recent advancements in deep reinforcement learning (DRL) has markedly broadened their applications in real-world robotic tasks, including locomotion control [8, 9], perception [10, 11], and decision making [12, 13].
Existing DRL-based exposure control methods [6, 7] suffer the online interative training scheme. Tomasi et al. [6] utilized a convolutional neural network (CNN) to dynamically adjust camera gain and exposure time settings. They devised a self-supervised labeling technique using dual cameras and a specialized capture pattern. However, their approach of online data collection combined with offline agent training presents complexities and inefficiencies in implementation. Additionally, the infrequent policy updates within their system may hinder the generalization capabilities of the agent. Lee et al. [7] proposed a device-specific method to facilitate online training. They vectorized input images to eliminate CNN modules, thereby reducing model complexity. They introduced a progressive lighting curriculum that incrementally exposed the agent to increasingly complex training scenarios. Nonetheless, the dependence on a specialized stable device limits their method’s generalizability and fails to address camera motion. Moreover, their approach to image vectorization sacrifices essential high-level visual information, such as feature detection and tracking.
@@ III-B Our System Design
In this work, we employ DRL to train agents for camera exposure control. Different from existing research [6, 7] which directly output the dual control parameters from the agent network, we split the exposure control problem into two distinct modules. The first module, referred as $\mathbf{ModuleExpo}$, determines the optimal exposure $e$ for the coming image, formulated as
@@ IV-A Photometric Synthesis Process
Inspired from Gamache et al. [17], we leverage bracketing image techniques to encompass a wide range of irradiances present within a scene. Specifically, we employ a bracketing capture consisting of $5$ images (Fig. 3(a)) with exposure times of $50$ us, $200$ us, $1$ ms, $5$ ms, and $20$ ms. When an agent requires an image with a specific target exposure, we select the seed image whose exposure time is just below the target, and then serves it as the basis for synthesizing, thus enabling effective interaction.
@@ IV-B Motion Augmentation
We realize the necessity of incorporating a broad spectrum of motion dynamics within our training datasets. However, the training trajectories remain invariant after collection. Additionally, employing a bracketing technique to widen the range of irradiance inherently diminishes the simulator’s frame rate output. This reduction in frame rates indirectly affects the perceived speed of camera movement, thus introducing discrepancies between training and inference.
@@ V-C3 Pose Error Reward
Sequence Built-in Shim [1] Zhang [2] Kim [4] Zhang [5] DRL-stat DRL-feat DRL-pose
Courtyard $\times$1 1.42 / 0.29 2.02 / 0.31 1.97 / 0.32 2.23 / 0.31 2.55 / 0.31 2.17 / 0.31 1.77 / 0.33 2.04 / 0.35
$\times$2 - (1) 7.98 / 0.88 6.78 / 0.71 2.45 / 0.53 1.64 / 0.49 1.50 / 0.49 1.75 / 0.52 2.38 / 0.58
Corridor $\times$1 1.98 / 0.22 2.24 / 0.24 2.14 / 0.25 2.26 / 0.25 2.19 / 0.24 1.96 / 0.23 0.82 / 0.21 3.00 / 0.32
$\times$2 4.28 / 0.39 (1) 5.26 / 0.44 (1) 5.72 / 0.46 (1) 6.12 / 0.56 (2) 4.88 / 0.45 (1) 4.13 / 0.43 (1) 1.29 / 0.32 7.48 / 0.64 (2)
Parking $\times$1 2.40 / 0.43 3.14 / 0.43 3.54 / 0.43 - (1) 1.44 / 0.42 2.34 / 0.42 1.98 / 0.41 2.69 / 0.51
$\times$2 - (1) - (1) - (1) - (3) - (2) - (3) 1.84 / 0.62 - (1)
Switch $\times$1 2.52 / 0.34 (1) 3.38 / 0.35 (1) 3.32 / 0.35 (1) 2.02 / 0.30 (1) 1.81 / 0.27 (1) 2.30 / 0.32 (1) 2.67 / 0.34 (1) 1.65 / 0.23
@@ VI-B Experiment: Visual Odometry
Several methods are implemented as the baseline comparison. Built-in is a feedback control method from FLIR cameras, aimed at optimizing average image brightness to a moderate value. Shim [1] and Zhang [2] are two gradient-based methods target on optimizing image gradient magnitude and its variants. Kim [4] and Zhang [5] are two function-fitting methods characterized by rapid parameter update speed. Since we designed the three reward functions, we trained DRL-stat, DRL-feat, and DRL-pose for comparison.
Quantitative results are presented in Table II. In the Courtyard sequence, performance across all methods is comparable with minimal variance. However, when operating at double speed, the Built-in was unable to adjust adequately to rapid motion, resulting in over-exposure and subsequent tracking failures. In the sequences of Corridor and Parking, characterized by significant lighting variations, DRL-feat demonstrates superior performance. This advantage stems from its capability to swiftly predict motion-based events and rapidly adjust to scene parameters, thereby maintaining consistent interframe tracking even at double speed. In contrast, other methods fail to deliver rapid and stable control under conditions of swift motion coupled with dramatic lighting changes, resulting in degradation or failure of the VO systems. In the Switch sequence, characterized by unpredictable lighting changes of $4$ times, all methods were interrupted by the most dramatic transition from dark to light, expect DRL-pose. This is because DRL-pose prefers a lower exposure: DRL-pose aims to directly minimize inter-frame pose errors. This enhances the agent’s preference for stable structural features over unstable textural features, which are plentiful but less reliable. Low exposure settings help obscure textural features and prevent blooming effects around structural features, which are typically delineated by contrasts between light and dark areas.
Additionally, we conducted a static experiment in the Switch scene to compare the response speeds of different methods to abrupt unpredictable lighting changes. The response performance is illustrated in Fig. 5. The function fitting method exhibited the quickest response, benefitting from its one-shot mechanism. Our DRL-based methods all respond faster than the traditional feedback control methods because their update hyperparameters are adaptively determined by the neural network.
Figure 5: The experiment of the reaction speed with a switch case of light off then on. The method of Zhang [5] performs the fastest because of its principle of one step control. All our DRL methods (DRL-stat, DRL-feat, and DRL-pose) respond faster than feedback control methods (Built-in and Shim).
@@ VI-D Experiment: Time Consumption
Our training process does not necessitate any interaction with camera hardware or the real environment, rendering it significantly faster than previous approaches [6, 7]. The training durations for our three agents—DRL-stat, DRL-feat, and DRL-pose—were $3$ hours $13$ minutes, $9$ hours $35$ minutes, and $10$ hours $33$ minutes, respectively, all trained for $10000$ episodes. The average training time per frame was recorded at $2.3$ ms, $6.9$ ms, and $7.6$ ms, respectively. In comparison, the method developed by Lee [7] with an online interaction platform, required an average of $100$ ms per frame. Consequently, our system demonstrates significantly higher training efficiency.